# Word problems and embedding-obstructions in cellular automata groups on groups

### v2

@@ 7 The lamplighter group
Let $C$ be a finite nontrivial abelian group, let $A$ be a nontrivial finite alphabet, $n\geq 2$, and $L=C\wr\mathbb{Z}$. Then $\mathrm{Aut}(A^{G})$ has co-NEXPTIME-complete word problems.
@@
We will assume $C=\mathbb{Z}_{2}$, because all the cases are similar (or rather, $\mathbb{Z}_{2}$ is the hardest case, as it needs a preprocessing step due to insufficient branching). We will also assume that $A$ is of specific form, as in the general case one can argue as in the proof of Theorem 4.1, i.e. work with a set of unbordered words representing symbols from a larger alphabet.
@@
Now as the alphabet we take $A=\mathbb{Z}_{5}\times\mathbb{Z}_{6}\times\mathbb{Z}_{7}\times W\times\{0,1,2\}^{2}\times B\times D$, where $B=C^{2}$ and $C$ is the set of colors of Wang tiles, and $D$ is an auxiliary finite set.
The basis of the proof is the same as in [4]. The lamplighter group contains large grids. More specifically, geometrically this group is a horocyclic product of two full binary trees, which means that we can obtain it as a connected component of a direct product of such trees, where one tree is inverted, and we have to move up and down simultaneously in the two trees. (Thus when in one tree a node branches, in the other we move to a parent.) Now, a full binary tree of depth $n$ naturally has $2^{n}$ elements on its frontier, and we can identify them with the numbers $[0,2^{n}-1]$. In the direct product, there is then a natural copy of a square $[0,2^{n}-1]^{2}$. The square happens to be in a single connected component of the horocyclic product, which will allow us to use automorphisms to ensure computation on the massive grid obtained by connecting ($\mathbb{Z}^{2}$-)neighboring elements of the square.
It is useful to step into a subgroup before describing the grids formally. Let $G$ be an index-$2$ subgroup of $L$ isomorphic to $\mathbb{Z}^{2}_{2}\wr\mathbb{Z}$. Write $L^{\prime}$ for the natural copy $\mathbb{Z}_{2}\times\{0\}\wr\mathbb{Z}$ of $L$ in $G$. Let $t$ be the natural generator of $\mathbb{Z}$, and write $a=t$, $b=t\cdot(1,0)$, $c=t\cdot(0,1)$ where $(1,0),(0,1)$ are the natural generators of $\mathbb{Z}_{2}^{2}$. So $a,b$ are a standard set of generators of $L^{\prime}$. The point of this is that for each $g\in L^{\prime}$, $g+\langle c\rangle$ is a copy (coset) of $\mathbb{Z}$ where we can encode counters, and for distinct $g,g^{\prime}\in L^{\prime}$ these cosets are disjoint.
@@
Write also $f(\ell,m)=f_{1}(\ell,m)$ so $f_{g}(\ell,m)=gf(\ell,m)$. Here, words are $1$-indexed and recall that binary representations are written with least significant bit first, i.e. LSB. Sometimes, we call the image of this function the grid, and $g$ is the root of the grid.
Note that $f(\ell,m)$ for different $\ell,m$ commute as elements of $L^{\prime}$ (the total shift of such an element is 00, so they belong in the $\bigoplus_{\mathbb{Z}}\mathbb{Z}_{2}$-part of $\bigoplus_{\mathbb{Z}}\mathbb{Z}_{2}\rtimes\mathbb{Z}$), and form a copy of $\mathbb{Z}_{2}^{(2^{n})^{2}}$ inside it, whose action can be described (through the identifying map $f$) as swapping adjacent rectangles at different levels. We think of the grid as being drawn on a plane $\mathbb{R}^{2}$ with north, south, west and east as the four directions. The elements $a,b$ move “up” in this point of view, For a picture of a piece of this group (drawn as described here), and a highlighted grid, see [4].
The counters of $g\in L$ are the two numbers read in the two first components of the $\{0,1,2\}^{3}$-component of the alphabet, read in positions $g,gc,\ldots,gc^{n-1}$, and interpreted in binary (if there are $2$s present, the counters are undefined).
@@
  • the counters at $f_{g}(\ell,m)$ have value $\ell,m$ respectively;
@@
If we can write a formula for such an element in polynomial time, then we have a reduction from the problem of checking whether a length $2^{n}$-computation of $T$ accepts the given input of size $n$, to the co-word problem of $\mathrm{Aut}(A^{L})$. The problem being reduced is indeed NEXPTIME-complete since $T$ is a universal Turing machine, and because Wang tiles simulate the Turing machine explicitly, i.e. they directly code a spacetime diagram of the Turing machine with the (south-north) height of the tiling corresponding exactly to computation time, and (west-east) width exactly to space usage. Note also that as in the case of the PSPACE-hardness proof in Lemma A.3, while we allow only an exponential time rather than $2^{O(n^{d})}$ as in the definition of NEXPTIME, we can always produce a polynomial padding in the reduction.
As in previous hardness proofs, it is difficult to an element as described above directly, since the grid is not directly connected inside $L$, and we have to do all the checking outside the grid. Again, we will use dyadically accelerated funneling.
@@
Just like in the proof of Theorem 3.1 we read words over $(0,1)^{*}2^{*}$ (indices) written on $(hc^{0},hc^{1},\ldots,hc^{n-1})$ for $h\in\{a,b\}^{\leq n}$ in the first $\{0,1,2\}$-component, and construct an automorphism $\psi_{1}^{\prime}$ of polynomial norm in $n$ that performs a nontrivial permutation $\pi$ at position $g$ in the $\mathbb{Z}_{7}$-component, if the positions $g\{a,b\}^{n}$ encode a valid (west-east row of a) $W$ tiling.
The basic idea is the same, i.e. an inductive argument using dyadic intervals $I$ of depths, which are read on the indices. However, now instead of having Turing machine configurations rooted at the various $gh$ with $h\in\{a,b\}^{\leq n}$, we have in each node just a single $B$-symbol, which codes $C^{2}$ where $C$ is the set of colors of $W$. When going up the tree, from $(c,c^{\prime})$ we produce $(c,c^{\prime\prime})$ in the $a$-branch and $(c^{\prime\prime},c)$ in the $b$-branch. The interpretation of $(c,c^{\prime})$ in a node is that $c$ (resp. $c^{\prime}$) is the west (resp. east) color of the leftmost leaf eventually produced from that node. The effect of this is to make the west and east colors consistent on the leaves. On the final $n$th level (which is where the grid of interest is), we check that the colors $(c,c^{\prime})$ are the left and right color of the Wang tile present in that node (it does not matter what the $W$-components contain outside the grid).
We should also ensure at this point that the final row has an accepting configuration. This is easy to include this in the check in the topmost dyadic interval of size $2$.
In addition to synchronizing the colors, the indices are directly used to synchronize the first $\{0,1,2\}$-components of the counters (to the correct grid value, relative to $g$). In the second $\{0,1,2\}$, we check that the tree contains a constant counter value, so that the south-north coordinate stays fixed.
@@
Note that for this to work, it is important that the $\mathbb{Z}_{6}$-content is not read or affected by $\psi^{\prime}$ (which is the case); and that the $\mathbb{Z}_{7}$-cells on a valid grid are never affected by $\psi_{1}^{\prime}$. The latter holds because when $\psi_{1}^{\prime}$ acts nontrivially, counters on the grid contain binary numbers in the first counter, while any cell permuted by $\psi_{1}^{\prime}$ has has $2$ in its first $\{0,1,2\}$-component (being the root of a connecting tree, thus having index $2^{n}$ in the terminology of the proof of Theorem 3.1).
@@
By symmetrically using trees that connect the grid from “above”, we have elements $\psi_{2}$ that only act nontrivially at $g$, in the $\mathbb{Z}_{6}$-component, if the “column” $f_{g}(0,j)$ has consistent colors in the Wang tiles, in the south and north directions, and the counters are also are consistent.
In the argument giving $\psi_{1}$, the basis is that we can perform a controlled permutation on the $\mathbb{Z}_{7}$-component based on coordinates being good (though here we only implicitly defined the notion of goodness). These formulas are ultimately commutator formulas, and the goodness checks are ultimately of the form “perform a permutation $\pi^{\prime}$ if the node is good”. Thus, we can equally well use the map $\psi_{1}$ in place of the notion of goodness, but now using the $\mathbb{Z}_{6}$-components, and use the trees above the grid connecting columns from above, analogously to the definition of $\psi_{2}$. Then we can use partial shifts (as in the transition from $\psi_{1}^{\prime}$ to $\psi_{1}$) to move this permutation to the $\mathbb{Z}_{5}$-component of a node on the grid the grid. This gives us an automorphism $\psi_{3}$ that performs a nontrivial permutation in the $\mathbb{Z}_{5}$ component at $g$ if and only if every west-east row of the grid $f_{g}([0,2^{n}-1],[0,2^{n}-1])$ rooted at $g$ is consistent (individually). Symmetrically from $\psi_{2}$ we can build, analogously to the definition of $\psi_{1}$, an automorphism $\psi_{4}$ that acts nontrivially if and only if all the columns are valid.
Now the commutator $[\psi_{3},\psi_{4}]$ has word norm $n$ in $\mathrm{PAut}(A)$, and performs a nontrivial permutation at $g$ if and only if every west-east row and every south-north column of the grid $f_{g}([0,2^{n}-1],[0,2^{n}-1])$ has consistent tiling and consistent Wang tiles.
@@ 8 Upper bounds and polynomial splitting schemes
In this section we show the co-NP, PSPACE and co-NEXPTIME upper bounds. In the case of co-NP, we also show completeness in this section (as this requires only a single sentence).
8.1 The co-NP upper bound
@@
For an f.g. group $G$ with polynomial growth, the automorphism group of an SFT with NP language has co-NP word problem.
@@
Polynomial growth is equivalent to virtually nilpotent, by Gromov’s theorem. It is known that these groups have polynomial-time word problem. The word problem of any f.g. subgroup can be solved by guessing a witness (which fits in a polynomial-sized ball due to the growth rate). Checking the witness is in the SFT is in NP by assumption, and checking that the automorphism acts nontrivially on it can be done in polynomial time by applying the local rules. ∎
@@
For an f.g. group $G$ with polynomial growth, $\mathrm{Aut}(A^{G})$ always has co-NP-complete word problem.
@@ 8.2 The PSPACE upper bound
We first consider a full shift $A^{G}$, and explain how to extend this to FEP subshifts. It suffices to show that the co-word problem is in alternating polynomial time AP, since co-PSPACE = PSPACE = AP (by determinism and Lemma A.2). Let $F$ be a finite set of automorphisms, and let $r$ be a bound on their radii. We may assume $r=1$ by changing the generating set $S$ of $G$. Given a word $w$ of length $k$ over $F$, we now describe how to prove the nontriviality, equivalently that there exists a pattern $p\in A^{B_{k}}$ such that composing the local rules of the cellular automata $w$ in sequence and applying them to $p$ gives a symbol distinct from $p_{1_{G}}$ at $1_{G}$.
@@
We show that we can figure out in alternating polynomial time whether such a pattern exists. To do this, we guess a polynomial sized subset $C_{1}$ of $B_{k}$. Then we compute its (polynomial-sized) boundary $C_{1}S\setminus C_{1}$. Note that we can check in polynomial space, and thus alternating polynomial time, whether two elements of the boundary are connected, and thus we can figure out a polynomial-size set of representatives of the boundary (or we can use $C_{1}S\setminus C_{1}$ directly and redo much of the computation). [^2] There is a little subtlety here: In alternating polynomial time computations, we may use PSPACE computations in recursive calls, which means we have alternating access to PSPACE. It may seem that this implies that APTIME trivially contains APSPACE, contradicting the fact APSPACE = EXPTIME and EXPTIME ${=}$ PSPACE = APTIME is a major open problem. The point is that alternating access from a polynomial time computation to PSPACE does not mean we have access to APSPACE, as in the latter we can do exponentially long computation including alternations. We also guess a pattern $p\in A^{C_{1}\times[0,k]}$ at this point.
The point is that we want to now check for each boundary element $g$, whether, letting $D$ be the connected component of the Cayley graph of $B_{k}\setminus C$ containing $g$, we can find a pattern $q\in A^{D\times[0,k]}$ which agrees with $p$ in $C\times[0,k]$, everywhere follows the local rules (in the sense described two paragraphs above), and that the contents of $1_{G}$ in $c_{k}$ differ from the contents of $1_{G}$ in $c_{0}$ if $1_{G}\in D$. The algorithm may not be correct for all $C,g$, [^3] In case we are not following a good sequence of splits, we may not shrink the set quickly enough, and the algorithm itself cannot easily check how much it shrinks the set at each point. but its positive claims are always going to be correct.
@@
Continuing this way, if a polynomial splitting scheme exists, in about $\log_{\alpha}|B_{k}|$ steps of recursion (which amounts to a linear amount of steps in $k$ since $\alpha>1$) we have figured out whether a witness exists: If one exists, we can use existential steps to ensure that the $C_{i}$ follow the polynomial splitting scheme, and that the patterns guessed on the $C_{i}$ come correspond to a pattern where $w$ acts nontrivially; when stepping to connected components, we use universal steps to ensure that all of the components admit such valid contents. On the other hand, the algorithm can only succeed if it manages to both find a polynomial splitting scheme, and a valid configuration where $w$ acts nontrivially.
@@
For a free group, the Cayley graph is a tree, and as the cuts one can use successive elements of geodesics: at each step, we have a single branch of the ball rooted on some $g$, and stepping forward by one step in the unique geodesic from $1_{G}$ to $g$ splits the branch in three sets of equal size, which gives a polynomial splitting scheme for $\alpha=2/3$. ∎
@@
As discussed above, it suffices to consider $g\geq 2$, since nonamenability is preserved by commensurability. Nonamenable surface groups are quasi-isometric to the hyperbolic plane [11]. Specifically, it is classical that the universal cover of the surface is the hyperbolic plane $H^{2}$, and by Švarc-Milnor the fundamental group is quasi-isometric to it. On the other hand, the hyperbolic plane is well-known to be quasi-isometric to the pentagon model. ∎
@@
It suffices to show that these exist in the pentagon model. Furthermore, it is easy to reduce the problem to finding a polynomial splitting scheme for “one-sided balls” $P_{k}=\{(x,y)\;|\;y\in[0,k],x\in[0,2^{y}-1]\}$, i.e. vertices reachable from $(0,0)$ by only moves that increment the second coordinate. Namely, these sets sandwich contain balls of the pentagon model, of roughly the same radius.
@@
For nonamenable free groups, nonamenable surface groups and their products with virtually nilpotent groups, the automorphism group of any PEPB subshift has PSPACE-complete word problem.
@@
These groups have polynomial splitting schemes, so the word problem of the automorphism group of any PEPB subshift is in PSPACE by Theorem 8.2. By Theorem 3.1, for the f.g. subgroup $\mathrm{PAut}$ the problem is PSPACE-hard in these cases, since nonamenable free groups and surface groups have free submonoids. ∎

### v3

@@ 8 The lamplighter group
Let $C$ be a finite nontrivial abelian group, let $A$ be a nontrivial finite alphabet, $n\geq 2$, and $L=C\wr\mathbb{Z}$. Then $\mathrm{Aut}(A^{G})$ has a finitely-generated subgroup with co-NEXPTIME-hard word problem.
@@
We will assume $C=\mathbb{Z}_{2}$, because all the cases are similar (or rather, $\mathbb{Z}_{2}$ is the hardest case, as it needs a preprocessing step due to insufficient branching). We will also assume that $A$ is of specific form, as in the general case one can argue as in the proof of Theorem 4.6, i.e. work with a set of unbordered words representing symbols from a larger alphabet (observing that the lamplighter group has plenty of self-embeddings).
@@
Now as the alphabet we take $A=\mathbb{Z}_{5}\times\mathbb{Z}_{6}\times\mathbb{Z}_{7}\times W\times\{0,1,2\}^{2}\times B$, where $B=C^{3}$ and $C$ is the set of colors of Wang tiles.
The basis of the proof is the same as in [4]. The lamplighter group contains large grids. More specifically, geometrically this group is a horocyclic product of two full binary trees, which means that we can obtain it as a connected component of a direct product of such trees, where one tree is inverted, and we have to move up and down simultaneously in the two trees. (Thus when in one tree a node branches, in the other we move to a parent.)
Now, a full binary tree of depth $n$ naturally has $2^{n}$ elements on its frontier, and we can identify them with the numbers $[0,2^{n}-1]$. In the direct product, there is then a natural copy of a square $[0,2^{n}-1]^{2}$. The square happens to be in a single connected component of the horocyclic product, which will allow us to use automorphisms to ensure computation on the massive grid obtained by connecting ($\mathbb{Z}^{2}$-)neighboring elements of the square.
It is useful to step into a subgroup before describing the grids formally. Let $G$ be a subgroup of $L$ isomorphic to $\mathbb{Z}^{2}_{2}\wr\mathbb{Z}$. Write $L^{\prime}$ for the natural copy $\mathbb{Z}_{2}\times\{0\}\wr\mathbb{Z}$ of $L$ in $G$. Let $t$ be the natural generator of $\mathbb{Z}$ in $G$ (i.e. the “shift map”), and write $a=t$, $b=t\cdot(1,0)$, $c=t\cdot(0,1)$ where $(1,0),(0,1)$ are the natural generators of $\mathbb{Z}_{2}^{2}$. So $a,b$ are a standard set of generators of $L^{\prime}$, and $a,b,c$ one for $G$. The point of this is that for each $g\in L^{\prime}$, $g+\langle c\rangle$ is a copy (coset) of $\mathbb{Z}$ where we can encode indices, and for distinct $g,g^{\prime}\in L^{\prime}$ these cosets are disjoint. We will not need to consider a generating set for $L$, and could work directly in $G$ from now on, but we can equally well work in $L$ directly, independently in each left $G$-coset.
@@
Write also $f(\ell,m)=f_{1}(\ell,m)$ so $f_{g}(\ell,m)=gf(\ell,m)$. Here, words are $1$-indexed and binary representations are written with least significant bit first. Sometimes, we call the image of this function the grid, and $g$ is the root of the grid.
Note that $f(\ell,m)$ for different $\ell,m$ commute as elements of $L^{\prime}$ (the total shift of such an element is 00, so they belong in the $\bigoplus_{\mathbb{Z}}\mathbb{Z}_{2}$-part of $G=\mathbb{Z}_{2}^{2}\wr\mathbb{Z}=\bigoplus_{\mathbb{Z}}\mathbb{Z}_{2}^{2}\rtimes\mathbb{Z}$), and form a copy of $\mathbb{Z}_{2}^{(2^{n})^{2}}$ inside it, whose action can be described (through the identifying map $f$) as swapping adjacent rectangles at different levels. Note that in particular that if $h$ is in the grid of $g$, then the grids of $g$ and $h$ agree as sets, but define different “orientations” for it.
When a root is fixed, we think of the grid as being drawn on a horizontal plane $\mathbb{R}^{2}$ with north, south, west and east as the four directions. The elements $a,b$ move “up” (out of the plane) in this point of view, For a picture of a piece of this group (drawn as described here), and a highlighted grid, see [4].
The grid coordinates of $g\in L$ is the pair of numbers read in the two first components of the $\{0,1,2\}^{2}$-component of the alphabet, read in positions $g,gc,\ldots,gc^{n-1}$, and interpreted in binary (if there are $2$s present, the grid coordinates are undefined).
@@
  • the grid coordinates of $f_{g}(\ell,m)$ are $(\ell,m)$;
@@
If we can write a formula for such an element in polynomial time, then we have a reduction from the problem of checking whether a length $2^{n}$-computation of $T$ accepts the given input of size $n$, to the co-word problem of $\mathrm{Aut}(A^{L})$. The problem being reduced is indeed NEXPTIME-complete since $T$ is a universal Turing machine, and because Wang tiles simulate the Turing machine explicitly, i.e. they directly code a spacetime diagram of the Turing machine with the (south-north) height of the tiling corresponding exactly to computation time, and (west-east) width exactly to space usage. Note also that (as in the case of the PSPACE-hardness proof in Lemma B.4) while we allow only an exponential time rather than $2^{O(n^{d})}$ as in the definition of NEXPTIME, we can always produce a polynomial padding in the reduction.
It is difficult to construct an element as described above directly, since the grid is not directly connected inside $L$, and we have to do all the checking outside the grid. Thus, besides cheching the above conditions on the grid, we have auxiliary requirements on connecting trees. Again, we will use ripple catching.
@@
Just like in the proof of Theorem 5.2 we read words over $(0,1)^{*}2^{*}$ (playing the role of the index) written on $(hc^{0},hc^{1},\ldots,hc^{n-1})$ for $h\in\{a,b\}^{\leq n}$ in the first $\{0,1,2\}$-component, and construct an automorphism $\psi_{1}^{\prime}$ of polynomial norm in $n$ that performs a nontrivial permutation $\pi$ at position $g$ in the $\mathbb{Z}_{7}$-component, if the positions $g\{a,b\}^{n}$ encode a valid (west-east row of a) $W$ tiling.
To implement this with ripple catching, we firstly include in the notion of successfulness of nodes that the index is incremented as in the proof of Theorem 5.2, i.e. at the root the index is $2^{n}$, and when moving up the tree, the generators are appended at the first $2$-symbol. In the second $\{0,1,2\}$-component of the alphabet, as we go up the tree we check that the values stay the same throughout the tree (and $2$ does not appear). All in all, once we land on the grid, the index is form $\{0,1\}^{n}$ and the second components also form a word in $\{0,1\}^{n}$, thus we have valid grid coordinates.
For the verification of correctness of the tiling, we have in each node of the tree just a single $B$-symbol, which codes $C^{3}$ where $C$ is the set of colors of $W$. When going up the tree, if at some node we have the triple $(c,c^{\prime\prime},c^{\prime})$, then in the successor relations we verify that we have $(c,c^{\prime\prime\prime},c^{\prime\prime})$ in the $a$-branch and $(c^{\prime\prime},c^{\prime\prime\prime\prime},c)$ in the $b$-branch, for some choices of $c^{\prime\prime\prime},c^{\prime\prime\prime\prime}$.
The interpretation of $(c,c^{\prime\prime},c^{\prime})$ in a node is that $c$ (resp. $c^{\prime}$) is the west (resp. east) color of the leftmost leaf eventually produced from that node. The role of the middle color $c^{\prime\prime}$ is technical: in our definition of successfulness through relations, we cannot directly synchronize the same $c^{\prime\prime}$ for the $a$- and $b$-successor (of course, we could allow this by slightly changing the formalism).
The effect of this is to make the west and east colors consistent on the leaves. On the final $n$th level (which is where the grid of interest is), we check that the colors $c,c^{\prime}$ of the triple $(c,c^{\prime\prime},c^{\prime})$ are the left and right color of the Wang tile present in that node (it does not matter what the $W$-components contain outside the grid, or what the $c^{\prime\prime}$-color is).
We should also ensure at this point that the final row has an accepting configuration. This is easy to include in the check performed in the nodes that land on grid coordinates.
@@
Note that for this to work, it is important that the $\mathbb{Z}_{6}$-content is not read or affected by $\psi^{\prime}$ (which is the case); and that the $\mathbb{Z}_{7}$-cells on a valid grid are never affected by $\psi_{1}^{\prime}$. The latter holds because when $\psi_{1}^{\prime}$ acts nontrivially, counters on the grid contain binary numbers in the first counter, while any cell permuted by $\psi_{1}^{\prime}$ has has $2$ in its first $\{0,1,2\}$-component (being the root of a connecting tree, thus having index $2^{n}$).
@@
By symmetrically using trees that connect the grid from “above”, we have elements $\psi_{2}$ that only act nontrivially at $g$, in the $\mathbb{Z}_{6}$-component, if the “column” $f_{g}(0,j)$ has consistent colors in the Wang tiles, in the south and north directions, and the counters are also consistent.
In the argument giving $\psi_{1}$, the basis is that we can perform a controlled permutation on the $\mathbb{Z}_{7}$-component based on coordinates being successful (though here we only implicitly defined the notion of successfulness). These formulas are ultimately commutator formulas, and the successfulness checks are ultimately of the form “perform a permutation $\pi^{\prime}$ if the node is successful”. Thus, we can equally well use the map $\psi_{1}$ in place of the notion of successfulness, but now using the $\mathbb{Z}_{6}$-components, and use the trees above the grid connecting columns from above, analogously to the definition of $\psi_{2}$.
Then we can use partial shifts (as in the transition from $\psi_{1}^{\prime}$ to $\psi_{1}$) to move this permutation to the $\mathbb{Z}_{5}$-component of a node on the grid. This gives us an automorphism $\psi_{3}$ that performs a nontrivial permutation in the $\mathbb{Z}_{5}$ component at $g$ if and only if every west-east row of the grid $f_{g}([0,2^{n}-1],[0,2^{n}-1])$ rooted at $g$ is consistent (individually). Symmetrically from $\psi_{2}$ we can build, analogously to the definition of $\psi_{1}$, an automorphism $\psi_{4}$ that acts nontrivially if and only if all the columns are valid.
Now the commutator $[\psi_{3},\psi_{4}]$ has word norm polynomial in $n$ in $\mathrm{PAut}(A)$, and performs a nontrivial permutation at $g$ if and only if every west-east row and every south-north column of the grid $f_{g}([0,2^{n}-1],[0,2^{n}-1])$ has consistent tiling and consistent Wang tiles.
@@ 9 Upper bounds and polynomial splitting schemes
In this section we show the co-NP, PSPACE and co-NEXPTIME upper bounds. In the case of co-NP, although it does not thematically fit, completeness is also shown in this section; the reason is that the proof requires only a single sentence.
9.1 The co-NP upper bound (and lower bound)
@@
For an f.g. group $G$ with polynomial growth, the automorphism group of an subshift with NP language has co-NP word problem.
@@
Polynomial growth is equivalent to virtually nilpotent, by Gromov’s theorem. It is known that these groups have polynomial-time word problem. The word problem of any f.g. subgroup can be solved by guessing a witness (which fits in a polynomial-sized ball due to the growth rate). Checking the witness is in the language of the subshift is in NP by assumption, and checking that the automorphism acts nontrivially on it can be done in polynomial time by applying the local rules. ∎
@@
For an f.g. infinite group $G$ with polynomial growth, $\mathrm{Aut}(A^{G})$ has a finitely-generated subgroup with co-NP-complete word problem.
@@ 9.2 The PSPACE upper bound
We first consider a full shift $A^{G}$, and then explain how to extend this to PEPB subshifts. It suffices to show that the co-word problem is in alternating polynomial time AP, since co-PSPACE = PSPACE = AP (the first equality is trivial from the determinism of the PSPACE model, and the second is Lemma B.3).
Let $F$ be a finite set of automorphisms, and let $r$ be a bound on their radii. We may assume $r=1$ by changing the generating set $S$ of $G$. Given a word $w$ of length $k$ over $F$, we now describe how to prove the nontriviality, equivalently that there exists a pattern $p\in A^{B_{k}}$ such that composing the local rules of the cellular automata $w$ in sequence and applying them to $p$ gives a symbol distinct from $p_{1_{G}}$ at $1_{G}$.
@@
We show that we can figure out in alternating polynomial time whether such a pattern exists. To do this, we guess a polynomial sized subset $C_{1}$ of $B_{k}$. Then we compute its (polynomial-sized) boundary $C_{1}S\setminus C_{1}$. Note that we can check in polynomial space, and thus alternating polynomial time, whether two elements of the boundary are connected, and thus we can figure out a polynomial-size set of representatives of the boundary (or we can use $C_{1}S\setminus C_{1}$ directly and redo much of the computation). [^4] There is a little subtlety here: In alternating polynomial time computations, we may use PSPACE computations in recursive calls, which means we have alternating access to PSPACE. It may seem that this implies that APTIME trivially contains APSPACE, which would be surprising since APSPACE = EXPTIME and EXPTIME ${=}$ PSPACE = APTIME is a major open problem. The point is that alternating access from a polynomial time computation to PSPACE does not mean we have access to APSPACE, as in the latter we can do exponentially long computation including alternations, only “around” these recursive calls. We also guess a pattern $p\in A^{C_{1}\times[0,k]}$ at this point.
The point is that we want to now check for each boundary element $g$, whether, letting $D$ be the connected component of the Cayley graph of $B_{k}\setminus C_{1}$ containing $g$, we can find a pattern $q\in A^{D\times[0,k]}$ which agrees with $p$ in $C_{1}\times[0,k]$, everywhere follows the local rules (in the sense described two paragraphs above), and that the contents of $1_{G}$ in $c_{k}$ differ from the contents of $1_{G}$ in $c_{0}$ if $1_{G}\in D$. The algorithm may not be correct for all $C_{1},g$, [^5] In case we are not following a good sequence of splits, we may not shrink the set quickly enough, and the algorithm itself cannot easily check how much it shrinks the set at each point. but its positive claims are always going to be correct.
@@
Continuing this way, if a polynomial splitting scheme exists, in about $\log_{\alpha}|B_{k}|$ steps of recursion (which amounts to a linear amount of steps in $k$ since $\alpha>1$) we have figured out whether a witness exists: If one exists, we can use existential steps to ensure that the $C_{i}$ follow the polynomial splitting scheme, and that the patterns guessed on the $C_{i}$ come correspond to a pattern where $w$ acts nontrivially; when stepping to connected components, we use universal steps to ensure that all of the components admit such valid contents. On the other hand, the algorithm only makes correct positive claims whether or not polynomial splitting schemes exist.
@@
Note that it is not enough in the above proof to assume that the subshift has PSPACE word problem, since we do not at any point directly check inclusion in the language, but only look locally for occurrences of finitely many forbidden patterns.
@@
For a free group, the Cayley graph is a tree, and as the cuts one can use successive elements of geodesics: at each step, we have a single branch of the ball rooted on some $g$, and stepping forward by one step in the unique geodesic from $1_{G}$ to $g$ splits the branch in three sets of equal size, which gives a polynomial splitting scheme with $\alpha=2/3$. ∎
@@
Thus, it suffices to consider the genus $2$ surface group with presentation $\langle a_{1},b_{1},a_{2},b_{2}\;|\;[a_{1},b_{1}][a_{2},b_{2}]\rangle$.
@@
Nonamenable surface groups are quasi-isometric to the hyperbolic plane [13]. Specifically, it is classical that the universal cover of the surface is the hyperbolic plane $H^{2}$, and by Švarc-Milnor the fundamental group is quasi-isometric to it. On the other hand, the hyperbolic plane is well-known to be quasi-isometric to the pentagon model. ∎
@@
It suffices to show that these exist in the pentagon model. Furthermore, it is easy to reduce the problem to finding a polynomial splitting scheme for “one-sided balls” $P_{k}=\{(x,y)\;|\;y\in[0,k],x\in[0,2^{y}-1]\}$, i.e. vertices reachable from $(0,0)$ by only moves that increment the second coordinate. Namely, these sets contain balls of the pentagon model, of roughly the same radius.
@@
For free groups, surface groups and their products with virtually nilpotent groups, the automorphism group of any PEPB subshift has PSPACE word problem. Furthermore, in the nonamenable case, automorphism groups of full shifts have f.g. subgroups whose word problems are PSPACE hard.
@@
These groups have polynomial splitting schemes, so the word problem of the automorphism group of any PEPB subshift is in PSPACE by Theorem 9.4. The second claim follows from Theorem 5.2, since nonamenable free groups and surface groups have free submonoids. ∎